# Advanced Encryption Standard AI artificial intelligence

### v1

@@ 5.2 FPGA Threats & Defenses
We showcased the feasibility of reverse engineering a real-world FPGA design, even within the limitations of a resource-constrained academic setting. Compared to regular ICs, the entry bar to reverse engineering is much lower. Unlike FPGAs, for which reliable open-source bitstream documentation projects are readily available [15, 32], the extraction of a netlist from an IC requires complex tooling, expensive equipment, and specialized expertise for preparation, imaging, and image analysis [3, 4, 5, 6, 7]. Given the inherent challenges of IC reverse engineering, we assume that FPGAs are vulnerable to reverse engineering by less powerful adversaries. In contrast, ICs that are more likely to be targeted by sophisticated attackers with nation-state level resources. On top of that, extracting an error-free netlist is much more likely for FPGAs than it is for ICs, since for the latter, error-prone sample preparation and image analysis is needed while an error-free bitstream documentation is more straight-forward to achieve. A correctly recovered gate-level netlist enables techniques like virtual probing (and simulation in general) that are sensitive to errors in the analyzed netlist, thereby increasing the threat potential even further.
In addition to IP theft, large-scale reverse engineering efforts can also provide an entry point for hardware Trojan insertion given the reconfigurability of FPGAs [48, 49, 50]. Consequently, a natural defense against netlist extraction would be to not use FPGAs for sensitive IP altogether. We recognize that this is not a viable option in many cases, hence additional FPGA countermeasures should be considered to defend against IP theft and Trojan insertion.
@@
A first line of defense is to impede extraction of a gate-level netlist. Proprietary bitstream formats may initially raise the bar for a reverse engineering attack, but given increasingly automated efforts, this may be overcome within reasonable time. This is inevitable since the user will always be able to translate HDL designs into bitstreams using EDA tools. As offered by major FPGA vendors, reliable cryptographic schemes can ensure bitstream confidentiality, integrity, and authentication. By encrypting and signing a bitstream, attacks such as IP theft are impeded. However, such bitstream protections are not always provided on low-cost FPGAs and legacy devices, which may continue to be in use for decades.
@@
In light of recent attacks on bitstream encryption schemes [21, 16, 17, 18, 19, 20], further precautions must be taken to harden the netlist against static and dynamic analysis by deteriorating those netlist properties that facilitate automated reverse engineering. Effective netlist obfuscation must aim at preventing the recovery of word-level structures as these structures are crucial for further analysis [51]. Dynamic reverse engineering is key to achieve an algorithmic understanding of larger circuits. Hence, to harden against such analysis techniques, non-simulatable primitives such as physical unclonable functions [52] or partial reconfiguration [38, 53, 38] should be included in the design and merged into the control and data paths.
5.3 Related Work
@@
Existing netlist reverse engineering techniques often only tackle isolated problems and have been evaluated on (sometimes outdated) open-source benchmarks rather than black-box designs. Hence, a comprehensive real-world case study is still lacking from literature. For a summary of prior work on netlist reverse engineering, see Azriel et al. [23].
Revisiting Related Work
@@
One research strand focuses on automated state registers recognition and FSM recovery [39, 35, 54, 55, 34]. RELIC by Meade et al. [37, 34] rates the similarity of FF fan-in trees to determine state and non-state FFs. Chowdhury et al. [56] propose a graph neural network (GNN) base approach that can also separate control FFs from data. To identify known sub-circuits, functional [57, 58, 59, 34] and graph-similarity based approaches [60] have been proposed. WordRev by Li et al. [58] comprises methods to recover word-level structures for functional matching. Subramanyan et al. [59] builds on their work to identify components such as adders, multipliers, counters, and registers. Albartus et al. introduced DANA [42] to identify high-level registers in a gate-level netlist. Meade et al. presented REBUS and REWIND to recover the datapath using similarity metrics comparable to RELIC [34]. Alrahis et al. [61] utilize GNN to determine for each gate whether it belonged to a module that implements known functionality. Narayanan et al. [47] also propose an identification of arithmetic operations by taking advantage of carry chains on FPGAs, as discussed in Section 4.2. Some work has been done to partition a netlist based on the inter connectivity inside a cluster of gates. Werner et al. [62] evaluate the applicability of conventional graph clustering algorithms for clustering in circuits, while Hong et al. [63] search for partitions by optimizing a n-cut with a GNN.
Advocating for Open-Source Implementations
In the software community, there is an agreement on common (open-source) tools and methods. This is not the case for hardware reverse engineering. Existing netlist reverse engineering techniques often lack reference implementations, posing challenges to reproducibility. During our case study, we encountered difficulties in applying existing techniques due to the effort required for re-implementation, surpassing the scope of our work.
We acknowledge the complexity of this issue, understanding that providing open-source implementations may clash with industry and government interests, occasionally even triggering concerns related to export regulations. We have personally encountered similar challenges. Despite these considerations, we advocate for a more open approach to contribute to the collective progress in netlist reverse engineering research, enabling the FPGA community to tackle more complex case studies, such as entire SoCs, based on our insights. This transparency is crucial for fostering a comprehensive understanding of the associated threats. We assume that such efforts are mostly conducted by governments behind closed doors, often away from public scrutiny.
5.4 Limitations and Future Work
@@
Our case study, while focused on a single DSP design on a relatively small FPGA, still provides valuable insights. As we have shown in Section 4, our case study findings inform generalized techniques that are applicable to implementations across different architectures, designs, and sizes. Future research could build up on this work to explore more elaborate real-world designs on more complex FPGAs to further underline the threat potential of IP theft for FPGAs. Our work builds the foundation for such case studies in that it provides the algorithms and implementations to dissect such a netlist at scale. Also, we have shown that there still is a need for better automated techniques for control logic extraction and analysis in gate-level netlists.
6 Conclusion
@@
Our work demonstrates the challenges associated with IP theft through FPGA reverse engineering and discusses the challenges faced in a real-world setting. To this end, our case study on a Lattice iCE40 FPGA that is part of iPhone 7 underscores the vulnerability of FPGAs to IP theft due to the low barrier for netlist extraction compared to regular ICs. Thereby, we reveal that while IP theft from FPGAs requires substantial effort, specialized skills, and domain knowledge, some of the most labor-intensive steps can be automated. We contribute to the field by introducing generalized netlist reverse engineering techniques and open-source implementations, which reduce manual effort and facilitate future research in this domain. Our techniques prove effective across different FPGA architectures and highlight the threat of IP theft even by less resourceful adversaries. This work emphasizes the importance of robust bitstream and netlist protections to safeguard valuable IP in FPGAs. Finally, we call for more open and transparent research in FPGA reverse engineering to better understand and mitigate associated threats.
Acknowledgments
@@
We thank Marc Fyrbiak and Max Hoffmann for the fruitful discussions. Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany´s Excellence Strategy - EXC 2092 CASA - 390781972, through ERC grant 695022, NSF grants CNS-1563829 and CNS-1749845, and ISF project IZ25-5793-2019-43 (“CHIoSec”).
@@ Appendix A Simulation in HAL
For netlist simulation in HAL, e.g., in the context of virtual probing, we extended the framework with an interface for commercial-grade simulation tools such as Verilator [64]. Furthermore, we added a waveform viewer to the HAL graphical user interface (GUI) that can be controlled via a Python application programming interface (API), see Figure 13. The dynamic state of the netlist can also be represented by coloring the nets in the graph view according to their current signal value, which adapts whenever the cursor in the waveform viewer is moved. See Figure 11 and Figure 12 for pictures of our actual virtual probing setup as initially described in Figure 7.
Figure 11: Side view of our custom breakout board sitting between the iPhone PCB and Maggie.
Figure 12: Top view of a Saleae logic analyzer being connected to the iPhone via our custom breakout board.
Figure 13: Screenshot of HAL showing the module tree on the left, the graph view in the middle, and the simulator on the right. The nets are colored depending on their current value, blue represents a 0 and red a 1.
@@ Appendix B Details on the Maggie DSP
Maggie’s data path takes two input signals supplied via Inter-IC Sound (I2S) and Serial Peripheral Interface (SPI) from the motor driver and Homer, respectively. These two signals are combined (scaled and subtracted) and then passed to a look-up table-based function where linear interpolation is used to enhance resolution. We believe that this step translates sensor measurements to the current LRA position, relying on calibration.
@@
The input processing result and the closed-loop controller output are passed to a processing stage which implements a state observer, estimating the LRA’s position and velocity. In particular, Maggie appears to implement a proportional-integral (PI) observer [65].
@@
The closed-loop controller takes a time series of desired actuator positions as its reference signal. Thus, as different haptic events are to be generated, this signal is changed. For instance, it is updated with the haptic feedback setting visible in Figure 12. Using the reference signal and the state observer outputs, the closed-loop controller accumulates the position and the velocity error (the desired velocity is obtained from numerical differentiation of the reference signal) and subtracts the integrated position estimation error. The particular combination of error components suggests that the control loop is a proportional-integral-derivative (PID)-controller [66, 67, 68]. Before being passed to the motor driver via I2S, a low-pass filter is applied to the output of the closed-loop controller.
Appendix C MUX Evaluation
For the evaluation of our word-level MUX detection using DANA, refer to Figure 14. The figure illustrates the five most frequently recovered MUX sizes for each benchmark, once synthesized for a Xilinx 7-series FPGA and once for s Lattice iCE40 FPGA. In this analysis, we exclusively consider MUXes comprising at least four gates. For the hilbert_transformer, no word-level MUXes were identified, hence we omitted it in Figure 14. We also provide our results for Maggie, which are only available for Lattice.
@@ Appendix C MUX Evaluation
Figure 14: Evaluation of our word-level MUX detection using DANA.

### v2

@@ 5.2. FPGA Threats & Defenses
We showcased the feasibility of reverse engineering a real-world FPGA design, even within the limitations of a resource-constrained academic setting. Compared to regular ICs, the entry bar to reverse engineering is much lower. Unlike FPGAs, for which reliable open-source bitstream documentation projects are readily available (Alliance, 2023b, a), the extraction of a netlist from an IC requires complex tooling, expensive equipment, and specialized expertise for preparation, imaging, and image analysis (Lippmann et al., 2019; Fyrbiak et al., 2017; Quadir et al., 2016; Torrance and James, 2011, 2009). Given these inherent challenges of IC reverse engineering, we assume that ICs are more likely to be targeted by sophisticated attackers with nation-state-level resources. In contrast, FPGAs are vulnerable to reverse engineering by less powerful adversaries. On top of that, extracting an error-free netlist is much easier for FPGAs than it is for ICs, since for the latter, error-prone sample preparation and image analysis is needed while error-free bitstream documentation is more straight-forward to achieve. A correctly recovered gate-level netlist enables techniques like virtual probing (and simulation in general) that are sensitive to errors in the netlist, thereby increasing the threat potential even further.
In addition to IP theft, large-scale reverse engineering efforts can also provide an entry point for hardware Trojan insertion given the reconfigurability of FPGAs (Kataria et al., 2019; Swierczynski et al., 2017, 2015a). Consequently, a natural defense against netlist extraction would be to not use FPGAs for sensitive IP altogether. We recognize that this is not a viable option in many cases, hence additional FPGA countermeasures should be considered to defend against IP theft and Trojan insertion.
@@
A first line of defense is to impede the extraction of a gate-level netlist. Proprietary bitstream formats may initially raise the bar for a reverse engineering attack, but given increasing automation, this may be overcome within a reasonable time. This is inevitable since the user will always be able to translate HDL designs into bitstreams using EDA tools. As offered by many FPGA vendors, robust cryptographic schemes can ensure bitstream confidentiality, integrity, and authentication. By encrypting and signing a bitstream, attacks such as IP theft are impeded. However, such bitstream protections are not always provided on low-cost FPGAs and legacy devices, which continue to be used for decades.
@@
In light of recent attacks on bitstream encryption schemes (Ender et al., 2020; Moradi et al., 2011, 2012; Moradi and Schneider, 2016; Swierczynski et al., 2015b; Tajik et al., 2017), further precautions must be taken to harden the netlist against static and dynamic analysis by deteriorating those netlist properties that facilitate automated reverse engineering. Effective netlist obfuscation must aim at preventing the recovery of word-level structures as these structures are crucial for further analysis (Basiashvili et al., 2022). Dynamic reverse engineering is key to achieve an algorithmic understanding of larger circuits. Hence, to harden against such analysis techniques, non-simulatable primitives such as physical unclonable functions (Wendt and Potkonjak, 2014) or partial reconfiguration (Fyrbiak et al., 2018; Stolz et al., 2021; Fyrbiak et al., 2018) should be included in the design and merged into the control and data paths.
5.3. Related Work
@@
Existing netlist reverse engineering techniques often only tackle isolated problems and have been evaluated on (sometimes outdated) open-source benchmarks rather than black-box designs. Hence, a comprehensive real-world case study is still lacking in the literature. For a summary of prior work, see Azriel et al. (Azriel et al., 2021).
@@
One research strand focuses on automated state registers recognition and FSM recovery (Shi et al., 2010; Meade et al., 2016b; McElvain, 2001; Meade et al., 2016a, 2018a). RELIC by Meade et al. (Brunner et al., 2019; Meade et al., 2018a) rates the similarity of FF fan-in trees to determine state and non-state FFs. Chowdhury et al. (Chowdhury et al., 2021) propose a graph neural network (GNN) base approach that can also separate control FFs from data. To identify known sub-circuits, functional (Li et al., 2012, 2013; Subramanyan et al., 2013; Meade et al., 2018a) and graph-similarity based approaches (Fyrbiak et al., 2020) have been proposed. WordRev by Li et al. (Li et al., 2013) comprises methods to recover word-level structures for functional matching. Subramanyan et al. (Subramanyan et al., 2013) builds on their work to identify components such as adders, multipliers, counters, and registers. Albartus et al. introduced DANA (Albartus et al., 2020) to identify high-level registers in a gate-level netlist. Meade et al. presented REBUS and REWIND to recover the datapath using similarity metrics comparable to RELIC (Meade et al., 2018a). Alrahis et al. (Alrahis et al., 2022) utilize GNN to determine for each gate whether it belonged to a module that implements known functionality. Narayanan et al. (Narayanan et al., 2023) also propose an identification of arithmetic operations by taking advantage of carry chains on FPGAs, as discussed in Section 4.2. Some work has been done to partition a netlist based on the interconnectivity of a cluster of gates. Werner et al. (Werner et al., 2018) evaluate the applicability of conventional graph clustering algorithms to circuits, while Hong et al. (Hong et al., 2023) search for partitions by optimizing an n-cut with a GNN.
5.4. Limitations and Future Work
@@
Our case study, while focused on a single DSP design on a relatively small FPGA, still provides valuable insights. As we have shown in Section 4, our case study findings inform generalized techniques that are applicable to implementations across different architectures, designs, and sizes. Future research could build up on this work to explore more elaborate real-world designs on more complex FPGAs to further underline the threat potential of IP theft for FPGAs. Our work builds the foundation for such case studies by providing the techniques to dissect such a netlist at scale. Also, we have shown that there still is a need for better automated techniques for control logic extraction and analysis in gate-level netlists.
6. Conclusion
@@
Our work highlights the challenges associated with IP theft through FPGA reverse engineering and discusses the issues arising in a real-world setting. To this end, our case study on a Lattice iCE40 FPGA that is part of iPhone 7 underscores the vulnerability of FPGAs to IP theft due to the low barrier for netlist extraction compared to regular ICs. Thereby, we reveal that while IP theft from FPGAs requires substantial effort, specialized skills, and domain knowledge, most labor-intensive steps can be automated. We contribute to the field by introducing generalized netlist reverse engineering techniques and open-source implementations, which reduce manual effort and facilitate future research in this domain. Our techniques prove effective across different FPGA architectures and highlight the threat of IP theft even by less resourceful adversaries. This work emphasizes the importance of robust bitstream and netlist protections to safeguard valuable IP in FPGAs. Finally, we call for more open and transparent research in FPGA reverse engineering to better understand and mitigate associated threats.
Acknowledgements.
@@
We thank Marc Fyrbiak and Max Hoffmann for the fruitful discussions. Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany´s Excellence Strategy - EXC 2092 CASA - 390781972, through ERC grant 695022, NSF grants CNS-1563829 and CNS-1749845, and ISF project IZ25-5793-2019-43 (CHIoSec).
@@ Appendix A Simulation in HAL
For netlist simulation in HAL, e.g., in the context of virtual probing, we extended the framework with an interface for commercial-grade simulation tools such as Verilator (Snyder, 2003). Furthermore, we added a waveform viewer to the HAL GUI that can be controlled via a Python API, see Figure 13. The dynamic state of the netlist can also be represented by coloring the nets in the graph view according to their current signal value, which adapts whenever the cursor in the waveform viewer is moved. See Figure 11 and Figure 12 for pictures of our actual virtual probing setup as initially described in Figure 7.
Figure 11. Side view of our custom breakout board sitting between the iPhone PCB and Maggie.
Figure 12. Top view of a Saleae logic analyzer being connected to the iPhone via our custom breakout board.
Figure 13. Screenshot of HAL showing the module tree on the left, the graph view in the middle, and the simulator on the right. The nets are colored depending on their current value, blue represents a 0 and red a 1.
@@ Appendix B Details on the Maggie DSP
Maggie’s data path takes two input signals supplied via I2S and SPI from the motor driver and Homer, respectively. These two signals are combined (scaled and subtracted) and then passed to a look-up table-based function where linear interpolation is used to enhance resolution. We believe that this step translates sensor measurements to the current LRA position, relying on calibration.
@@
The input processing result and the closed-loop controller output are passed to a processing stage which implements a state observer, estimating the LRA’s position and velocity. In particular, Maggie appears to implement a proportional-integral (PI) observer (Beale and Shafai, 1989).
@@
The closed-loop controller takes a time series of desired actuator positions as its reference signal. Thus, as different haptic events are to be generated, this signal is changed. For instance, it is updated with the haptic feedback setting visible in Figure 12. Using the reference signal and the state observer outputs, the closed-loop controller accumulates the position and the velocity error (the desired velocity is obtained from numerical differentiation of the reference signal) and subtracts the integrated position estimation error. The particular combination of error components suggests that the control loop is a proportional-integral-derivative (PID)-controller (Ellis, 2002, 2012; Wu et al., 2017). Before being passed to the motor driver via I2S, a low-pass filter is applied to the output of the closed-loop controller.
Appendix C Benchmarks
Table 3 depicts the resource utilization of our six benchmark designs. Differences in resource utilization between the Xilinx and Lattice benchmarks, e.g., for fft64, can be the result of the Xilinx synthesizer using DSPs where the Lattice synthesizer resorts to carry chains. Furthermore, Xilinx FPGAs leverage 6-input LUTs while Lattice FPGAs use 4-input ones.
Table 3. Open-source benchmark resource utilization in the number of gates.
Design Vendor LUT Carry FF BRAM DSP Other Total
ibex (lowrisc, [n. d.]) Xilinx 3503 65 1937 0 1 727 6233
Lattice 6379 295 1963 16 0 78 8731
icicle (grahamedgecombe, [n. d.]) Xilinx 1525 65 1009 270 0 78 2947
Lattice 2705 262 983 20 0 59 4029
simple_risc_v (damdoy, [n. d.]) Xilinx 1617 76 1003 6 0 37 2739
Lattice 2876 274 1235 6 0 31 4422
canny_edge_detector (angelobacchini OpenCores, [n. d.]) Xilinx 2538 636 3713 24 0 72 6983
Lattice 3298 2170 4484 32 0 68 10052
fft64 (Sergiyenko and Uzenkov, [n. d.]) Xilinx 1722 212 1988 3 4 160 4089
Lattice 5362 2008 2556 14 0 124 10064
sha256 (secworks, [n. d.]) Xilinx 2196 104 1830 0 0 269 4399
Lattice 3935 346 2378 0 0 90 6749
hilbert_transformer (OpenCores, [n. d.]) Xilinx 304 68 587 0 0 85 1044
Lattice 445 225 569 2 0 78 1349
maggie Lattice 3241 422 1342 18 4 19 5046
Appendix D MUX Evaluation
MUXes of sizes that divide the bit-size of the datapath or are multiples of a byte are likely to be correct. To this end, we plotted the distributions of the top 5 MUX sizes for all benchmarks in Figure 14. The figure illustrates the five most frequently recovered MUX sizes for each benchmark, once synthesized for a Xilinx 7-series FPGA and once for s Lattice iCE40 FPGA. In this analysis, we exclusively consider MUXes comprising at least four gates. For the hilbert_transformer, no word-level MUXes were identified, hence we omitted it in Figure 14. We also provide our results for Maggie, which are only available for Lattice.
@@ Appendix D MUX Evaluation
Figure 14. Evaluation of our word-level MUX detection using DANA.
Appendix E White-Box Case Study Comparison
In our white-box case study, we used an open-source Hilbert transformer design (OpenCores, [n. d.]) as the ground truth to validate our generalized techniques. A signal processing expert unaware of the nature of the analyzed design inspected the Python code representing the recovered Boolean functions to draw a block diagram. From their analysis, they correctly identified the Hilbert transformer. Figure 15 shows the block diagrams representing the implemented signal processing chain. The diagram recovered from the gate-level netlist is shown in 15(a) and the one taken from the implementation’s documentation (Kumm, 2008) in 15(b). While the recovered block diagram largely resembles the one from the documentation, partially matching exactly, there are also notable differences. Some are due to visualization choices, e.g., in favor of symmetry in the lower signal path, and others are due to implementation details affecting intermediate signals. For example, delay elements (represented by negative powers of $z$) can easily be moved around by factoring them into or out of parts of the processing. Similarly, delay elements that simply shift the entire output signal may be added or removed without altering the time-invariant processing behavior. Furthermore, the trivial addition with 00 denoted in the documentation (first dashed box labeled $C_{+90}(z)$ in 15(b)) is not found in the netlist, likely due to synthesizer optimizations. Please note that, despite the deviation of the visual representations and minor trivial differences, both block diagrams still describe the same overall processing.
(a) Block diagram based on recovered Boolean functions.
(b) Block diagram from official documentation (OpenCores, [n. d.]; Kumm, 2008).
Figure 15. Comparison of Hilbert transformer block diagrams in the white-box case study.